OpenStudy (anonymous):

Find x if f(x) = 2x + 7 and f(x) = -1.

OpenStudy (yellowlegoguy99):

First Multiply 2 x -1 = -2, then add -2 by 7, and you get 5. So x = 5.

OpenStudy (yellowlegoguy99):

Did that help?

OpenStudy (anonymous):

Yess, thank youu!

OpenStudy (yellowlegoguy99):

Could you give me a medal? :)

OpenStudy (anonymous):

actually, you get 2x+7=-1 2x=-7-1 2x=-8 x=-4

OpenStudy (yellowlegoguy99):

@thegeolance I believe you scored incorrectly

OpenStudy (anonymous):

x=-4